Solve the following system by the addition method.
-18x+42y= -10
Y= 3/7x

Answer by stanbon(57424)

Solve the following system by the addition method.
-18x+42y= -10
Y= 3/7x
----
Rearrange the 2nd equation:
7y = 3x
Then 3x - 7y = 0
-------
System:
-18x + 42y = -10
3x - 7y = 0
-----
Multiply thru the lower equation by 6:
------
System:
-18x + 42y = -10
18x - 42y = 0
-----
Adding you get:
0 = -10
-----
The system is inconsistent.
There is no solution for the system.
==============
Cheers,
Stan H.